FINANCIAL ACCOUNTING Chapter Inventory (3 multiple choice questions)

3)On March 1, Miles Ltd. had $24402 in inventory. During the month, the company purchased inventory worth $191467. An inventory count on March 31 revealed inventory on hand of $20399. Assuming Miles uses a periodic inventory system, what is the cost of goods available for sale for the month of March?
Select one:
a. $215869
b. $195470
c. $24402
d. $236268

answer: b?

4)Mika Enterprises uses a perpetual inventory system. The company had the following inventory and purchases for the month of March:

Mar. 1
Beginning Inventory
35 @ $19.82
Mar. 8
Purchase
103 @ $25.19
Mar. 16
Sale
44 @ $29.82

Calculate the cost of goods sold for the year assuming the company uses a FIFO cost flow system.

Select one:
a. $1312
b. $1108
c. $920
d. $2368

answer: c?

5)An entity has the following inventory transactions for an item of inventory:

Opening Inventory
113 units @ 10
$1130

Purchase 1
212 units @ 13
$2756

Sale 1
57 units

What is cost of goods sold assuming the entity uses the periodic weighted average method?
Select one:
a. $570
b. $741
c. $682
d. $656
Answer #1

3) The formula here applied is,

Beginning Inventory + Purchases - COGS = Ending Inventory (on hand)

putting the values from the questions,

24402 + 191467 - COGS = 20399

Therefore COGS = 24402 + 191467 - 20399 = 195470.

Answer is B.

4) In FIFO we dispose off the inventory which was from the beginning and when the First's are sold then we come to the purchased inventory.

here, our beginning inventory is 35 units @ 19.82

and purchased inventory is 103 units @ 25.19

units sold is 44 units. Here, first we will deduct the beginning inventory first the we will come to purchased inventory.

so, COGS = 35*19.82 + 9 (44-35)(remaining units) * 25.19

= 920.

The answer is C.

5) In weighted avg method, we take the weighted averages of the inventory available to us.

therefore, the weighted avg of units available for sale = (113*10 + 212*13)/325(total no. of units) = 11.957

COGS = 11.957 * 57(no. of units to be sold) = 682.

The answer is C.